OpenStudy (anonymous):

Given the function y = 5x – 7, how does the y-value change as x increases by 1?

OpenStudy (cj7529):

Sub in x+1 instead of x y=5(x+1)-7 y=5x+5-7 y=5x-7 +5 so y will increase by 5 when x increases by 1